### v1

# Disentangled World Models: Learning to Transfer Semantic Knowledge from Distracting Videos for Reinforcement Learning

Abstract
Training visual reinforcement learning (RL) in practical scenarios presents a significant challenge, i.e., RL agents suffer from low sample efficiency in environments with variations. While various approaches have attempted to alleviate this issue by disentanglement representation learning, these methods usually start learning from scratch without prior knowledge of the world. This paper, in contrast, tries to learn and understand underlying semantic variations from distracting videos via offline-to-online latent distillation and flexible disentanglement constraints. To enable effective cross-domain semantic knowledge transfer, we introduce an interpretable model-based RL framework, dubbed Disentangled World Models (DisWM). Specifically, we pretrain the action-free video prediction model offline with disentanglement regularization to extract semantic knowledge from distracting videos. The disentanglement capability of the pretrained model is then transferred to the world model through latent distillation. For finetuning in the online environment, we exploit the knowledge from the pretrained model and introduce a disentanglement constraint to the world model. During the adaptation phase, the incorporation of actions and rewards from online environment interactions enriches the diversity of the data, which in turn strengthens the disentangled representation learning. Experimental results validate the superiority of our approach on various benchmarks.
1 Introduction
Visual reinforcement learning (VRL) presents a promising approach for training agents within complex environments [14, 23, 8, 38, 16]. However, VRL frequently suffers from performance degradation in practical scenarios due to the complexity, volatility, and visual distractions in environments. Even minor environmental variations can result in significant pixel-level shifts, making the trained VRL policies ineffective or suboptimal [4, 3]. For instance, a slight change in lighting conditions can affect an object’s appearance (e.g., color, shadow, or other visual attributes). Therefore, it is crucial to enhance models with interpretability, enabling them to perceive, learn, and understand the semantic environmental variations.
Figure 1: Overview of our proposed framework. The key idea is to leverage distracting videos for semantic knowledge transfer, enabling the downstream agent to improve sample efficiency on unseen tasks.
Table 1: MuJoCo (downstream domain) vs. DMC (accessible distracting videos).
Video: DMC Target: MuJoCo Similarity / Difference
Task Reacher Easy Pusher Relevant robotic control tasks
Dynamics Two-link planar Multi-jointed robot arm Different
Action space Box(-1, 1, (2,), float32) Box(-2, 2, (7,), float32) Different
Reward range [0, 1] [-4.49, 0] Different
Disentangled representation learning (DRL) presents a promising approach to addressing the interpretability challenges inherent in the “black-box” nature of deep learning algorithms. Fundamentally, DRL approaches mimic the cognitive processes of biological intelligence, where understanding the world is facilitated by decomposing observations into distinct and independent factors  [2, 10, 34, 32, 35]. In this form, when a factor of variation is changed (e.g., color), only a small portion of features in the disentangled representation will be affected, enabling the agent to recover performance quickly. Several studies have explored the integration of DRL algorithms in the domain of VRL. For example, Higgins et al. [11] trained a $\beta$-VAE offline to obtain disentangled representations for reinforcement learning. TED [4] adopts a self-supervised auxiliary task to learn temporally disentangled representations for reinforcement learning. Additionally, Dunion et al. [3] introduced conditional mutual information to achieve disentangled representation that on top of conditional independence. However, existing methods typically learn the representations from scratch, lacking any prior knowledge of the world. These approaches often require extensive interactions with the environment to acquire desired behaviors.
Towards this challenge, we introduce a model-based interpretable VRL framework, dubbed Disentangled World Models (DisWM), which leverages prior knowledge extracted from distracting videos to facilitate the learning of unseen downstream tasks through latent distillation. Note that inspired by Distracting DeepMindControl [26], we build videos with visual distractions (e.g., color variations), which called distracting videos. Specifically, as depicted in Figure 1, our framework consists of two phases: first, we pretrain a DRL encoder to learn disentangled latent representations from distracting videos. By doing so, the pretrained DRL encoder is “knowledgeable” in terms of representation disentanglement. Subsequently, we finetune an orthogonally designed world model with dual constraints of disentanglement and distillation, leveraging semantic knowledge transferred from the pretrained model via offline-to-online latent distillation. Another benefit of disentangled world model adaptation is that incorporating actions and rewards from online interactions with the environment enriches the diversity of the visual observations, which in turn strengthens the process of disentangled representation learning. It is worth mentioning that, as a cross-domain framework, DisWM does not require the pretraining videos to originate from the same domain as the downstream tasks.
Experimental results demonstrate the effectiveness of our proposed approach in improving the sample efficiency of VRL agents across our modified DeepMind Control and MuJoCo Pusher. The contributions of this work can be summarized as follows:
  • We frame the problem of learning interpretable VRL agents as a domain transfer learning problem. The key idea is to extract semantic knowledge from distracting videos and transfers this disentanglement capability to downstream control tasks.
  • We present DisWM, an approach that follows the pretraining-finetuning paradigm using distracting videos, incorporating specific techniques of offline-to-online latent distillation and flexible disentanglement constraints.
2 Problem Setup
We formulate visual reinforcement learning as a partially observable Markov decision process (POMDP) that using DMC and MuJoCO Pusher as the test bench. Specifically, we concentrate on scenarios where videos without actions and rewards are accessible, enabling world knowledge transfer. The goal is to maximize the cumulative reward of the target POMDP $\left\langle\mathcal{O},\mathcal{A},\mathcal{T},\mathcal{R},\gamma\right\rangle$ by transferring the shared world knowledge from the videos. These notations correspond to the visual observation space, the action space, the transition probabilities, the reward function, and the discount factor, respectively.
For instance, in one of the cross-domain experiments, we use MuJoCo as the downstream domain and the frames collected from DMC as the distracting video. Table 1 highlights the differences between the two domains in terms of visual appearances, physical dynamics, action spaces, and reward functions.
3 Method
Figure 2: Architecture of Disentangled World Models. Left: The action-free video prediction model with disentanglement constraints is pretrained on distracting videos offline for the well-disentangled latent variable $\textbf{z}_{\text{disen}}$, which extracts semantic knowledge from the visual observations. The disentangled capability of $\textbf{z}_{\text{disen}}$ is then transferred to the world model through latent distillation. Right:  The action-conditioned world model is finetuned through disentanglement regularization online, which encourages the factorized representation. Moreover, the incorporation of actions and rewards enriches the diversity of the data, which in turn strengthens the disentangled representation learning.
3.1 Overview of DisWM
In this section, we present the details of DisWM, which involves three main stages (see Figure 2):
  1. a)
    Disentangled representation pretraining: Pretrain a DRL-based video prediction model from distracting videos to extract disentangled features.
  2. b)
    Offline-to-online latent distillation: Transfer the semantic knowledge from the pretrained model to the world model via cross-domain latent distillation.
  3. c)
    Disentangled world model adaptation: Finetune the downstream agent with disentanglement constraints by incorporating the action and reward information.
3.2 Disentangled Representation Pretraining
[…]
$\beta$-VAE encoder: ${\mathbf{z}}_{t}$ $=e_{\phi^{\prime}}(o_{t})$ (1)
Posterior state: ${z}_{t}$ $\sim q_{\phi^{\prime}}(z_{t}\mid z_{t-1},\mathbf{z}_{t})$
Prior state: $\hat{z}_{t}$ $\sim p_{\phi^{\prime}}(\hat{z}_{t}\mid z_{t-1})$
Reconstruction: $\hat{o}_{t}$ $\sim p_{\phi^{\prime}}(\hat{o}_{t}\mid z_{t})$
Isotropic unit Gaussian: $p(\mathbf{z})$ $=\mathcal{N}(\mathbf{0},I).$
where $\phi^{\prime}$ denotes the parameters of the model. The $\beta$-VAE-based video prediction model is trained to minimize the following loss function:
(ϕ)=superscriptitalic-$\mathcal{L}(\phi^{\prime})=$ $\mathbb{E}_{q_{\phi^{\prime}}}\Big{[}\sum_{t=1}^{T}\underbrace{-\ln p_{\phi^{\prime}}(o_{t}\mid z_{t})}_{\text{image reconstruction}}$ (2)
$+\underbrace{\beta_{1}\mathrm{KL}\left[q_{\phi^{\prime}}\left(z_{t}\mid z_{t-1},o_{t}\right)\|p_{\phi^{\prime}}\left(\hat{z}_{t}\mid z_{t-1}\right)\right]}_{\text{action-free KL loss }}$
$\underbrace{+\beta_{2}\mathrm{KL}\left[q_{\phi^{\prime}}(\mathbf{z}_{t}\mid o_{t})\ \|\ p(\mathbf{z}_{t})\right]}_{\text{disentanglement loss}}\Big{]}.$
The variantional posterior distribution $q_{\phi^{\prime}}(\mathbf{z}_{t}\mid o_{t})$ is encouraged to be close to the standard multivariate Gaussian distribution $\mathcal{N}(\mathbf{0},I)$ to strengthen the orthogonality and disentanglement of the latent space. The importance of the disentanglement loss term is governed by $\beta_{2}$.
3.3 Offline-to-Online Latent Distillation
After the offline pretraining with the distracting videos, the model is fine-tuned online to adapt to the downstream task by integrating actions and rewards (Lines 13 of Alg. 1). A straightforward approach to transfer disentangled features involves is to initialize the action-conditioned world model with checkpoints obtained from the pretrained video prediction model. Nevertheless, it may experience a potential mismatch issue caused by the discrepancies between the two domains in visual appearances and physical dynamics. We observe that directly applying the pretraining-finetuning paradigm for downstream tasks tends to overwrite the disentangled information encoded in the pretrained latent features, leading to the decreased performance (see Figure 4).
Through comprehensive pretraining on distracting videos that contain diverse visual variations, the video prediction model thus builds an interpretable and orthogonality latent space. In this space, the latent variable $\mathbf{z}_{\text{disen}}$ achieves a high degree of disentanglement. To exploit the prior semantic knowledge from the pretrained model and improve the sample efficiency of the downstream tasks, we introduce an offline-to-online latent distillation. This approach enables the disentangling capability of $\mathbf{z}_{\text{disen}}$ from the pretrained model to be effectively transferred to the latent variable $\mathbf{z}_{\text{task}}$ of the world model. Specifically, this is achieved by minimizing the Kullback-Leibler (KL) divergence between the latent distributions of the two domains. The corresponding distillation loss $\mathcal{L}_{\text{distill}}$ can be formulated as follows:
$\mathcal{L}_{\text{distill}}=\text{KL}\left(\mathbf{z}_{\text{disen}}\|\mathbf{z}_{\text{task }}\right)=\sum\mathbf{z}_{\text{disen}}\cdot\log\left(\frac{\mathbf{z}_{\text{disen }}}{\mathbf{z}_{\text{task }}}\right)$ (3)
3.4 Disentangled World Model Adaptation
By obtaining well-disentangled representations of $\mathbf{z}_{\text{disen}}$ and employing the latent distillation for knowledge transfer, we then propose a DRL-based world model $\mathcal{M}_{\phi}$, designed to harness these features to enhance interoperability and robustness against environmental variations (Lines 14-15 of Alg. 1). The components of $\mathcal{M}_{\phi}$ can be detailed as follows:
Recurrent transition: $h_{t}$ $=f_{\phi}(h_{t-1},z_{t-1},a_{t-1})$ (4)
$\beta$-VAE encoder: ${\mathbf{z}}_{t}$ $\sim e_{\phi}(o_{t})$
Posterior state: ${z}_{t}$ $\sim q_{\phi}(z_{t}\mid h_{t},\mathbf{z}_{t})$
Prior state: ${\hat{z}}_{t}$ $\sim p_{\phi}(\hat{z}_{t}\mid h_{t})$
Reconstruction: $\hat{o}_{t}$ $\sim p_{\phi}(\hat{o}_{t}\mid h_{t},z_{t})$
Reward prediction: $\hat{r}_{t}$ $\sim r_{\phi}(\hat{r}_{t}\mid h_{t},z_{t})$
Discount factor: $\hat{\gamma}_{t}$ $\sim p_{\phi}(\hat{\gamma}_{t}\mid h_{t},z_{t})$
Isotropic unit Gaussian: $p(\mathbf{\mathbf{z}})$ $=\mathcal{N}(\mathbf{0},I),$
where $\phi$ represents the combined parameters of the world model. We train $\mathcal{M}_{\phi}$ on the sampled data from the replay buffer $\mathcal{B}$ with the following loss function:
(ϕ)=italic-$\mathcal{L}(\phi)=$ $\mathbb{E}_{q_{\phi}}\Big{[}\sum_{t=1}^{T}\underbrace{-\ln p_{\phi}(o_{t}\mid h_{t},z_{t})}_{\text{image reconstruction}}\underbrace{-\ln r_{\phi}(r_{t}\mid h_{t},z_{t})}_{\text{reward prediction}}$ (5)
$\underbrace{-\ln p_{\phi}(\gamma_{t}\mid h_{t},z_{t})}_{\text{discount prediction }}\underbrace{+\alpha\ \mathrm{KL}\left[q_{\phi}(z_{t}\mid h_{t},o_{t})\ \|\ p_{\phi}(\hat{z}_{t}\mid h_{t})\right]}_{\text{KL divergence}}$
$\underbrace{+\beta\mathrm{KL}\left[q_{\phi}(\mathbf{z}_{t}\mid o_{t})\ \|\ p(\mathbf{z}_{t})\right]}_{\text{disentanglement loss}}+\underbrace{\eta\mathcal{L}_{\text{distill}}}_{\text{distillation loss}}\Big{]}.$
where $\beta$ is a hyperparameter used to balance reconstruction quality and disentanglement capability. In this adaptation statge, $\eta$ serves as a hyperparameter that gradually decreases from 0.1 to 0.01. Intuitively, $\eta$ controls the progressive adaptation of the world model with the shared world knowledge transfer from the pretrained video prediction model.
Through the comprehensive training processes of this framework, we equip the DisWM with the capability to learn and understand the underlying semantic representations. This enhancement enables the model to be less sensitive by the environmental variations, such as changes in object colors, positions, and backgrounds. Furthermore, by incorporating actions and rewards during the finetuning phase, the world model can generate data with more diverse representations, thereby improving the disentangled representation learning. For behavior learning, we utilize the actor-critic method that is in line with DreamerV2 [7] (Lines 16-18 of Alg. 1). For more details of behavior learning, please refer to the supplementary material.
Algorithm 1 The training pipeline of DisWM.
1:Hyperparameters: $H$: Horizon of latent imagination.
2:Require: Distracting video dataset $\mathcal{D}$.
3:Initialize: Parameters of the model $\{\phi,\psi,\xi\}$.
4:for training step $t=1,2,\ldots,K_{1}$ do $\triangleright$ Disentagnled representation pretraining
5:    Sample random minibatch $\{o_{t}\}_{t=1}^{T}\sim\mathcal{D}.$
6:    Obtain Gaussian prior $\mathbf{z}_{t}$ from Isotropic unit Gaussian $\mathcal{N}(\mathbf{0},I)$.
7:    Pretrain the action-free video prediction model with disentanglement regularization by minimizing Eq. (2).
8:end for
9:Train the random agent and collect a replay buffer $\mathcal{B}$.
10:while not converged do
11:    for training step $t=1,2,\ldots,K_{2}$ do
12:         Sample $\{(o_{t},a_{t},r_{t})\}_{t=1}^{T}\sim\mathcal{B}$.
13:         Distill the disentangled features to the world model using Eq. (3). $\triangleright$ Offline-to-online latent distillation
14:         Obtain Gaussian prior $\mathbf{z}_{t}$ from Isotropic unit Gaussian $\mathcal{N}(\mathbf{0},I)$. $\triangleright$ Disentangled world model adaptation
15:         Train the world model $\mathcal{M}_{\phi}$ with latent distillation and disentanglement constraints using Eq. (5).
16:         Generate $\{(\hat{z}_{i},\hat{a}_{i})\}_{i=t}^{t+H}$ using $\pi_{\psi}$ and $\mathcal{M}_{\phi}$. $\triangleright$ Behavior learning
17:         Train the critic $v_{\xi}$ over $\{(\hat{z}_{i},\hat{a}_{i})\}_{i=t}^{t+H}$.
18:         Train the actor $\pi_{\psi}$ over $\{(\hat{z}_{i},\hat{a}_{i})\}_{i=t}^{t+H}$.
19:    end for
20:    $o_{1}$leftarrow env.reset() $\triangleright$ Environment interaction
21:    for time step $t=1,2,\ldots,T$ do
22:         Sample $\hat{a}_{t}$ $\sim$ $\pi_{\psi}(\hat{a}_{t}\mid\hat{z}_{t})$.
23:         $r_{t},o_{t+1}\leftarrow$ env.step($\hat{a}_{t}$).
24:    end for
25:    Append data to the replay buffer $\mathcal{B}$.
26:end while
4 Experiments
4.1 Experimental Setups
Benchmark
We evaluate DisWM on DeepMind Control Suite (DMC) [29] and MuJoCo Pusher [30]. DMC is a widely adopted benchmark with comprehensive and flexible robotic-control tasks. For DMC benchmark, we use 5 tasks, i.e., Walker Walk, Cheetah Run, Hopper Stand, Finger Spin, Cartpole Swingup. In MuJoCo Pusher, a multi-jointed robotic arm with is employed to manipulate a target cylinder (obejct). The goal is to move the object to a designated target position using the robot’s end effector (fingertip). The agent receives the negative reward, which is a combination of three components: the distance between the fingertip and the target, the distance between the object and the target goal position, and a penalty for large actions.
Compared methods.
We compare DisWM with strong visual RL agents, including
  • DreamerV2 [7]: A model-based RL (MBRL) approach that trains world model and learns by imagining future latent states.
  • APV [24]: It learns informational representations via action-free pretraining on videos and finetunes the agent with learned representations in the downstream tasks with action.
  • DV2 Finetune: It pretrains a DreamerV2 agent [7] on distracting videos and then finetunes the trained model in the downstream tasks. Note that some tasks have different action spaces, which makes it difficult to finetune directly. Therefore, the action space of two tasks is set as the maximum action space of both environments.
  • TED [4]: It adopts a classification task to learn temporally disentangled representations in visual RL.
  • CURL [14]: A model-free RL method that employs contrastive learning to improve its sample efficiency.
Implementation details.
The visual observations of online finetuning stages are resized to $64\times 64$ pixels. Inspired by APV [24], we build distracting video datasets with 1M frames using DreamerV2 [7] to interact with the environments with visual color distractors. This video datasets consist of the samples stored in the replay buffer throughout the training process until the agent achieves maximum score. For tasks in the DMC benchmark, the training steps of the agent is limited to $1\times 10^{6}$ environment steps. Each run of DisWM requires roughly $5$GB of VRAM and takes around $16$ hours to train on a single RTX 3090 GPU. The dimensions of well-disentangled latent $\mathbf{z}_{\text{disen}}$ and downstream task latent $\mathbf{z_{\text{task}}}$ are both set to 20 in our approach. In Figure 3, we showcase the example observations of various tasks with color distractors. We train the agent using a fixed set of colors, where the RGB values are varied within a restricted range around the original values. Furthermore, at the midpoint of the training process, we change to a different color scheme for varying distractors.
Figure 3: Example image observations of our modified DMC and MuJoCo Pusher with color distractors.
[…]
Figure 4: Comparison of DisWM against visual RL baselines, including DreamerV2 [7], APV [24], DV2 Finetune, TED [4], CURL [14].
Figure 5: Visualization of traversals of $\beta$-VAE during the pretraining phase.
Figure 6: Fine-grained disentanglement results on MuJoCo Pusher during world model adaptation phase. Each row in Pusher images displays the traversal results on a specific attribute.
4.2 Main Comparison.
We evaluate the sample efficiency and task performance for all the methods with the training curves of episode return. Figure 4 shows the performance of DisWM and all the baselines. Remarkably, it achieves better performance than TED [4], a method on top of RAD [13] and tailored for environments with distractors. For the offline-to-online finetuning models, DV2 Finetune achieves the second-best performance by transferring knowledge from the distracting videos. However, we observe a significant decline in sample efficiency, particularly in scenarios with large data distribution shifts between the source and target domains (e.g., DMC $\rightarrow$ MuJoCo). These shifts can occur in various aspects, including visual observation, physical dynamics, reward definition, or the action space of the robots. Another crucial baseline is APV [24], which focuses on transferring knowledge obtained from videos with a stacked latent prediction model. Nevertheless, without environment-specific designs for visual distractors, directly training may eventually results in a decrease in performance in the downstream tasks. The CURL model struggles to learn effective behavior policies, especially in Hopper Stand.
Additionally, we present qualitative results in Figure 5 and Figure 6. Figure 5 shows the traversals of $\beta$-VAE during the pretraining phase. In each row of travesals, a distinct attribute varies, while other attributes remain constant, indicating that the pretrained model has successfully disentangled and learned this attribute, thereby improving the sample efficiency of the RL agent. Figure 6 displays the fine-grained disentanglement results on MuJoCo Pusher during the finetuning phase, demonstrating that the proposed world model can effectively disentangle the variations.
4.3 Model Analyses
Ablation studies.
Figure 7: These figures illustrate the ablation studies and sensitivity analyses of DisWM on DMC Walker Walk $\rightarrow$ Cheetah Run. Left: Comparison with DisWM without cross-domain latent distillation or disentanglement constraints. Middle: The sensitivity analyses of cross-domain latent distillation weight. Right: The performance of DisWM with different disentanglement scale.
We conduct ablation studies to validate the effect of the latent distillation and disentanglement constraints. Figure 7  (Left) shows corresponding results in the DMC Walker Walk $\rightarrow$ Cheetah Run. The green curve shows that removing the latent distillation of DisWM results in a decreased performance, which indicates that the latent distillation is essential during the early training stage. For the model represented by the blue curve, we do not adopt disentanglement constraints for both pretrain and finetune stages. It can be seen that the necessity of introducing DRL-based training and disentangled representation significantly improves the learning efficiency of the agent.
Sensitivity analyses.
We conduct sensitivity analyses on DMC (Cheetah Run $\rightarrow$ Walker Walk). As shown in Figure 7 (Middle), we observe that when $\beta$ for the representation disentanglement is too small, the model learns entangled latent representations. When $\beta$ is too large, it will impede the reconstruction of image, leading to a decline in performance. Latent distillation weight $\eta$ controls the cross-domain transfer scale. Intuitively, setting this hyperparameter too low may result in the downstream agent not getting enough knowledge from the pretrained models. Conversely, excessively high $\eta$ may result in the model overfitting to the pretrained model, which is not conducive to the learning of the downstream task.
Effects of video domain.
In Figure 8, we evaluate DisWM by pretraining on alternative video datasets, including frames collected from Walker Walk, Reacher Easy, and Finger Spin. Interestingly, comparing with the original DreamerV2 agent without pretraining, DisWM can always benefit from pretraining via offline-to-online latent distillation. It obtains the semantic knowledge from the pretrained model and strengthens the disentanglement capability during finetuning phase.
Figure 8: Performance of DisWM with different video datasets.
5 Related Work
Visual MBRL.
Visual RL learns control policies from raw pixels, which has achieved remarkable performance in various tasks, while prior RL studies focus on learning policies from low-dimensional states. Extant approaches can be divided into two main directions: model-free RL [14, 37, 27, 36, 20, 40, 15] and model-based RL [5, 6, 7, 9, 39, 25, 19, 17, 1, 22, 31, 16]. The following methods specifically address the variations of environments in visual MBRL. Pan [23] et al. decompose visual dynamics into controllable and uncontrollable states through the optimization of inverse dynamics. SeeX [12] proposed a bi-level optimization framework that adpots separated world model and maximizes the task-relevant uncertainty. Orthogonal to these studies, our approach employs a DRL-based world model to alleviate the issue of visual variations.
Transfer RL
To facilitate the learning of unseen tasks, transfer RL [28, 41, 18, 20, 21, 31] leverages the knowledge learned from past tasks. One promising way is to transfer world knowledge from accessible videos to improve the downstream control. APV [24] established a pretraining-finetuning framework with a stacked latent prediction model and video-based intrinsic bonus. IPV [33] introduced contextualized world models that pretrained on diverse in-the-wild videos. It incorporates a context encoder that works alongside the latent dynamics model into the image encoder to capture rich contextual information. PreLAR [38] pretrained the world model with the derived meaningful actions from the action-free video using an inverse dynamics encoder. Different from these approaches, we propose a new solution to transfer world knowledge from distracting videos to improve the learning efficiency of downstream tasks via offline-to-online latent distillation.
6 Conclusions and Limitations
In this paper, we present a transfer RL dubbed DisWM, which addresses the challenge of environment variations in practical scenarios. Our key insight is to leverage the accessible distracting videos to facilitate the sample efficiency of downstream tasks to offer flexible disentanglement constraints. Specifically, we introduce disentangled representation pretraining, offline-to-online latent distillation, and disentangled world model adaptation to improve the downstream control. DisWM demonstrates superior performance than existing visual RL baselines across various benchmarks.
One limitation of our approach is that disentangled representation learning encounters challenges in complex environments. Exploring the non-stationary environments with more intricate variations, such as time-varying background video distractions, could further highlight the potential of our approach for practical scenarios.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grants 62302246, 62250062), the Natural Science Foundation of Zhejiang Province, China (Grant LQ23F010008), the IDT Foundation of Youth Doctoral Innovation (Grant S203.2.01.32.002), the Smart Grid National Science and Technology Major Project (Grant 2024ZD0801200), the Shanghai Municipal Science and Technology Major Project (Grant 2021SHZDZX0102), and the Fundamental Research Funds for the Central Universities. Additional support was provided by the project of Supporting Program for Young and Middle-aged Scientific and Technological Innovation Talents in Shenyang (Grant RC210488) and the project of Provincial Doctoral Research lnitiation Fund Program (Grant 2023-BS-214), the High Performance Computing Center at Eastern Institute of Technology, Ningbo, and Ningbo Institute of Digital Twin.

### v2

@@
 Training visual reinforcement learning (RL) in practical scenarios presents a significant challenge, i.e., RL agents suffer from low sample efficiency in environments with variations.
-While various approaches have attempted to alleviate this issue by disentanglement representation learning, these methods usually start learning from scratch without prior knowledge of the world.
+While various approaches have attempted to alleviate this issue by disentangled representation learning, these methods usually start learning from scratch without prior knowledge of the world.
 This paper, in contrast, tries to learn and understand underlying semantic variations from distracting videos via offline-to-online latent distillation and flexible disentanglement constraints.
@@
 Additionally, Dunion et al.
-[3] introduced conditional mutual information to achieve disentangled representation that on top of conditional independence.
+[6] introduced conditional mutual information to achieve a disentangled representation with correlated data.
 However, existing methods typically learn the representations from scratch, lacking any prior knowledge of the world.
@@
 Towards this challenge, we introduce a model-based interpretable VRL framework, dubbed Disentangled World Models (DisWM), which leverages prior knowledge extracted from distracting videos to facilitate the learning of unseen downstream tasks through latent distillation.
-Note that inspired by Distracting DeepMindControl [26], we build videos with visual distractions (e.g., color variations), which called distracting videos.
+It is crucial to note that distracting videos refer to videos with visual distractions, which are beneficial for learning disentangled representations.
 Specifically, as depicted in Figure 1, our framework consists of two phases: first, we pretrain a DRL encoder to learn disentangled latent representations from distracting videos.
@@
   • We frame the problem of learning interpretable VRL agents as a domain transfer learning problem.
-The key idea is to extract semantic knowledge from distracting videos and transfers this disentanglement capability to downstream control tasks.
+The key idea is to extract semantic knowledge from distracting videos and transfer this disentanglement capability to downstream control tasks.
   • We present DisWM, an approach that follows the pretraining-finetuning paradigm using distracting videos, incorporating specific techniques of offline-to-online latent distillation and flexible disentanglement constraints.
 2 Problem Setup
-We formulate visual reinforcement learning as a partially observable Markov decision process (POMDP) that using DMC and MuJoCO Pusher as the test bench.
+We formulate visual reinforcement learning as a partially observable Markov decision process (POMDP) that uses DMC and MuJoCo Pusher as the test bench.
 Specifically, we concentrate on scenarios where videos without actions and rewards are accessible, enabling world knowledge transfer.
@@
 Figure 2: Architecture of Disentangled World Models.
-Left: The action-free video prediction model with disentanglement constraints is pretrained on distracting videos offline for the well-disentangled latent variable $\textbf{z}_{\text{disen}}$, which extracts semantic knowledge from the visual observations.
+(a) The action-free video prediction model with disentanglement constraints is pretrained on distracting videos offline for the well-disentangled latent variable $\textbf{z}_{\text{disen}}$, which extracts semantic knowledge from the visual observations.
 The disentangled capability of $\textbf{z}_{\text{disen}}$ is then transferred to the world model through latent distillation.
-Right:  The action-conditioned world model is finetuned through disentanglement regularization online, which encourages the factorized representation.
+(b) The action-conditioned world model is finetuned through disentanglement regularization online, which encourages the factorized representation.
 Moreover, the incorporation of actions and rewards enriches the diversity of the data, which in turn strengthens the disentangled representation learning.
@@
 where $\phi^{\prime}$ denotes the parameters of the model. The $\beta$-VAE-based video prediction model is trained to minimize the following loss function:
-(ϕ)=superscriptitalic-$\mathcal{L}(\phi^{\prime})=$ $\mathbb{E}_{q_{\phi^{\prime}}}\Big{[}\sum_{t=1}^{T}\underbrace{-\ln p_{\phi^{\prime}}(o_{t}\mid z_{t})}_{\text{image reconstruction}}$ (2)
+$\mathcal{L}(\phi^{\prime})=$ $\mathbb{E}_{q_{\phi^{\prime}}}\Big{[}\sum_{t=1}^{T}\underbrace{-\ln p_{\phi^{\prime}}(o_{t}\mid z_{t})}_{\text{image reconstruction}}$ (2)
 $+\underbrace{\beta_{1}\mathrm{KL}\left[q_{\phi^{\prime}}\left(z_{t}\mid z_{t-1},o_{t}\right)\|p_{\phi^{\prime}}\left(\hat{z}_{t}\mid z_{t-1}\right)\right]}_{\text{action-free KL loss }}$
@@
 1).
-A straightforward approach to transfer disentangled features involves is to initialize the action-conditioned world model with checkpoints obtained from the pretrained video prediction model.
+A straightforward approach to transfer disentangled features involves initializing the action-conditioned world model with checkpoints obtained from the pretrained video prediction model.
 Nevertheless, it may experience a potential mismatch issue caused by the discrepancies between the two domains in visual appearances and physical dynamics.
-We observe that directly applying the pretraining-finetuning paradigm for downstream tasks tends to overwrite the disentangled information encoded in the pretrained latent features, leading to the decreased performance (see Figure 4).
+Directly applying the pretraining-finetuning paradigm for downstream tasks tends to overwrite the disentangled information encoded in the pretrained latent features, leading to decreased performance when there are large domain discrepancies between the source and the target domains.
 Through comprehensive pretraining on distracting videos that contain diverse visual variations, the video prediction model thus builds an interpretable and orthogonality latent space.
@@
 where $\phi$ represents the combined parameters of the world model. We train $\mathcal{M}_{\phi}$ on the sampled data from the replay buffer $\mathcal{B}$ with the following loss function:
-(ϕ)=italic-$\mathcal{L}(\phi)=$ $\mathbb{E}_{q_{\phi}}\Big{[}\sum_{t=1}^{T}\underbrace{-\ln p_{\phi}(o_{t}\mid h_{t},z_{t})}_{\text{image reconstruction}}\underbrace{-\ln r_{\phi}(r_{t}\mid h_{t},z_{t})}_{\text{reward prediction}}$ (5)
+$\mathcal{L}(\phi)=$ $\mathbb{E}_{q_{\phi}}\Big{[}\sum_{t=1}^{T}\underbrace{-\ln p_{\phi}(o_{t}\mid h_{t},z_{t})}_{\text{image reconstruction}}\underbrace{-\ln r_{\phi}(r_{t}\mid h_{t},z_{t})}_{\text{reward prediction}}$ (5)
 $\underbrace{-\ln p_{\phi}(\gamma_{t}\mid h_{t},z_{t})}_{\text{discount prediction }}\underbrace{+\alpha\ \mathrm{KL}\left[q_{\phi}(z_{t}\mid h_{t},o_{t})\ \|\ p_{\phi}(\hat{z}_{t}\mid h_{t})\right]}_{\text{KL divergence}}$
@@
 where $\beta$ is a hyperparameter used to balance reconstruction quality and disentanglement capability.
-In this adaptation statge, $\eta$ serves as a hyperparameter that gradually decreases from 0.1 to 0.01.
+In this adaptation stage, $\eta$ serves as a hyperparameter that gradually decreases from 0.1 to 0.01.
 Intuitively, $\eta$ controls the progressive adaptation of the world model with the shared world knowledge transfer from the pretrained video prediction model.
 Through the comprehensive training processes of this framework, we equip the DisWM with the capability to learn and understand the underlying semantic representations.
-This enhancement enables the model to be less sensitive by the environmental variations, such as changes in object colors, positions, and backgrounds.
+This enhancement enables the model to be less sensitive to environmental variations, such as changes in object colors, positions, and backgrounds.
 Furthermore, by incorporating actions and rewards during the finetuning phase, the world model can generate data with more diverse representations, thereby improving the disentangled representation learning.
@@
 1).
-For more details of behavior learning, please refer to the supplementary material.
+For more details of behavior learning, please refer to Supplementary Material B.
 Algorithm 1 The training pipeline of DisWM.
@@
 3:Initialize: Parameters of the model $\{\phi,\psi,\xi\}$.
-4:for training step $t=1,2,\ldots,K_{1}$ do $\triangleright$ Disentagnled representation pretraining
+4:for training step $t=1,2,\ldots,K_{1}$ do $\triangleright$ Disentangled representation pretraining
 5:  Sample random minibatch $\{o_{t}\}_{t=1}^{T}\sim\mathcal{D}.$
@@
 19:  end for
-20:    $o_{1}$leftarrow env.reset() $\triangleright$ Environment interaction
+20:  $o_{1}$ $\leftarrow$ env.reset() $\triangleright$ Environment interaction
 21:  for time step $t=1,2,\ldots,T$ do
@@
 Benchmark
-We evaluate DisWM on DeepMind Control Suite (DMC) [29] and MuJoCo Pusher [30].
+We evaluate DisWM on DeepMind Control Suite (DMC) [34], MuJoCo Pusher [35], and DrawerWorld [37].
 DMC is a widely adopted benchmark with comprehensive and flexible robotic-control tasks.
 For DMC benchmark, we use 5 tasks, i.e., Walker Walk, Cheetah Run, Hopper Stand, Finger Spin, Cartpole Swingup.
-In MuJoCo Pusher, a multi-jointed robotic arm with is employed to manipulate a target cylinder (obejct).
+In MuJoCo Pusher, a multi-jointed robotic arm is employed to manipulate a target cylinder (object).
 The goal is to move the object to a designated target position using the robot’s end effector (fingertip).
 The agent receives the negative reward, which is a combination of three components: the distance between the fingertip and the target, the distance between the object and the target goal position, and a penalty for large actions.
-Compared methods.
-We compare DisWM with strong visual RL agents, including
-  • DreamerV2 [7]: A model-based RL (MBRL) approach that trains world model and learns by imagining future latent states.
-  • APV [24]: It learns informational representations via action-free pretraining on videos and finetunes the agent with learned representations in the downstream tasks with action.
-  • DV2 Finetune: It pretrains a DreamerV2 agent [7] on distracting videos and then finetunes the trained model in the downstream tasks.
-Note that some tasks have different action spaces, which makes it difficult to finetune directly.
-Therefore, the action space of two tasks is set as the maximum action space of both environments.
-  • TED [4]: It adopts a classification task to learn temporally disentangled representations in visual RL.
-  • CURL [14]: A model-free RL method that employs contrastive learning to improve its sample efficiency.
+DrawerWorld is a modified Metaworld [43] benchmark designed to evaluate texture adaptability in manipulation tasks.
+It includes five additional textures from realistic photos and grid texture.
+During training, we initially employ the grid texture and then change to the wood texture midway, while adopting the metal texture exclusively for evaluation.
+The corresponding results on DrawerWorld are reported in Supplementary Material C.2.
+Furthermore, the introduction of compared baselines is given in Supplementary Material A.
 Implementation details.
@@
 This video datasets consist of the samples stored in the replay buffer throughout the training process until the agent achieves maximum score.
-For tasks in the DMC benchmark, the training steps of the agent is limited to $1\times 10^{6}$ environment steps.
-Each run of DisWM requires roughly $5$GB of VRAM and takes around $16$ hours to train on a single RTX 3090 GPU.
+For tasks in the DMC benchmark, the training steps of the agent are limited to $1\times 10^{6}$ environment steps.
+Each run of DisWM requires roughly 55GB of VRAM and takes around $16$ hours to train on a single RTX 3090 GPU.
 The dimensions of well-disentangled latent $\mathbf{z}_{\text{disen}}$ and downstream task latent $\mathbf{z_{\text{task}}}$ are both set to 20 in our approach.
@@
 Figure 6: Fine-grained disentanglement results on MuJoCo Pusher during world model adaptation phase. Each row in Pusher images displays the traversal results on a specific attribute.
-4.2 Main Comparison.
+4.2 Main Comparison
 We evaluate the sample efficiency and task performance for all the methods with the training curves of episode return.
@@
 Another crucial baseline is APV [29], which focuses on transferring knowledge obtained from videos with a stacked latent prediction model.
-Nevertheless, without environment-specific designs for visual distractors, directly training may eventually results in a decrease in performance in the downstream tasks.
+Nevertheless, without environment-specific designs for visual distractors, directly training may eventually result in a decrease in performance in the downstream tasks.
 The CURL model struggles to learn effective behavior policies, especially in Hopper Stand.
+Additional results on the challenging DMC Humanoid Walk can be found in Supplementary Material C.1.
 Additionally, we present qualitative results in Figure 5 and Figure 6.
 Figure 5 shows the traversals of $\beta$-VAE during the pretraining phase.
-In each row of travesals, a distinct attribute varies, while other attributes remain constant, indicating that the pretrained model has successfully disentangled and learned this attribute, thereby improving the sample efficiency of the RL agent.
-Figure 6 displays the fine-grained disentanglement results on MuJoCo Pusher during the finetuning phase, demonstrating that the proposed world model can effectively disentangle the variations.
+In each row of traversals, a distinct attribute varies, while other attributes remain constant, indicating that the pretrained model has successfully disentangled and learned this attribute, thereby improving the sample efficiency of the RL agent.
+Figure 6 displays the fine-grained disentanglement results on MuJoCo Pusher during the finetuning phase, demonstrating that the world model can effectively disentangle the variations.
 4.3 Model Analyses
@@
 As shown in Figure 7 (Middle), we observe that when $\beta$ for the representation disentanglement is too small, the model learns entangled latent representations.
-When $\beta$ is too large, it will impede the reconstruction of image, leading to a decline in performance.
+When $\beta$ is too large, it will impede the reconstruction of the image, leading to a decline in performance.
 Latent distillation weight $\eta$ controls the cross-domain transfer scale.
@@
 Conversely, excessively high $\eta$ may result in the model overfitting to the pretrained model, which is not conducive to the learning of the downstream task.
+Additional sensitivity analyses on the latent space dimension are provided in Supplementary Material C.3.
 Effects of video domain.
-In Figure 8, we evaluate DisWM by pretraining on alternative video datasets, including frames collected from Walker Walk, Reacher Easy, and Finger Spin.
-Interestingly, comparing with the original DreamerV2 agent without pretraining, DisWM can always benefit from pretraining via offline-to-online latent distillation.
+In Figure 8, we evaluate DisWM on DMC Cartpole Swingup by pretraining on alternative video datasets, including frames collected from Finger Spin, Reacher Easy, Walker Walk, and Hopper Stand.
+Interestingly, compared with the DreamerV2 agent without pretraining, DisWM can always benefit from pretraining via offline-to-online latent distillation.
 It obtains the semantic knowledge from the pretrained model and strengthens the disentanglement capability during finetuning phase.
-Figure 8: Performance of DisWM with different video datasets.
+Figure 8: Performance of DisWM on DMC Cartpole Swingup with different video datasets.
 5 Related Work
 Visual MBRL.
-Visual RL learns control policies from raw pixels, which has achieved remarkable performance in various tasks, while prior RL studies focus on learning policies from low-dimensional states.
+Visual RL learns control policies from raw pixels, which has achieved remarkable performance in various tasks [37, 3, 4, 31], while prior RL studies focus on learning policies from low-dimensional states.
 Extant approaches can be divided into two main directions: model-free RL [18, 44, 32, 42, 24, 47, 19] and model-based RL [8, 9, 10, 12, 46, 30, 23, 21, 1, 27, 36, 20].
@@
 decompose visual dynamics into controllable and uncontrollable states through the optimization of inverse dynamics.
-SeeX [12] proposed a bi-level optimization framework that adpots separated world model and maximizes the task-relevant uncertainty.
+SeeX [16] proposed a bi-level optimization framework that adopts a separated world model and maximizes the task-relevant uncertainty.
 Orthogonal to these studies, our approach employs a DRL-based world model to alleviate the issue of visual variations.
@@
 Exploring the non-stationary environments with more intricate variations, such as time-varying background video distractions, could further highlight the potential of our approach for practical scenarios.
-Acknowledgments
-This work was supported by the National Natural Science Foundation of China (Grants 62302246, 62250062), the Natural Science Foundation of Zhejiang Province, China (Grant LQ23F010008), the IDT Foundation of Youth Doctoral Innovation (Grant S203.2.01.32.002), the Smart Grid National Science and Technology Major Project (Grant 2024ZD0801200), the Shanghai Municipal Science and Technology Major Project (Grant 2021SHZDZX0102), and the Fundamental Research Funds for the Central Universities.
-Additional support was provided by the project of Supporting Program for Young and Middle-aged Scientific and Technological Innovation Talents in Shenyang (Grant RC210488) and the project of Provincial Doctoral Research lnitiation Fund Program (Grant 2023-BS-214), the High Performance Computing Center at Eastern Institute of Technology, Ningbo, and Ningbo Institute of Digital Twin.
+Acknowledgements.
+This work was supported by Grants of NSFC 62302246 & 62250062, ZJNSFC LQ23F010008, Ningbo 2023Z237 & 2024Z284 & 2024Z289 & 2023CX050011 & 2025Z038, the Smart Grid National Science and Technology Major Project (2024ZD0801200), the Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102), the Fundamental Research Funds for the Central Universities, and the IDT Foundation of Youth Doctoral Innovation (S203.2.01.32.002).
+Additional support was provided by the project of Supporting Program for Young and Middle-aged Scientific and Technological Innovation Talents in Shenyang (Grant RC210488), the project of Provincial Doctoral Research Initiation Fund Program (Grant 2023-BS-214), the High Performance Computing Center at Eastern Institute of Technology, Ningbo, and Ningbo Institute of Digital Twin.
+Supplementary Material
+Appendix A Compared Baselines
+We compare DisWM with strong visual RL agents, including
+  • DreamerV2 [10]: A model-based RL (MBRL) approach that trains world model and learns by imagining future latent states.
+  • APV [29]: It learns informational representations via action-free pretraining on videos and finetunes the agent with learned representations in the downstream tasks with action.
+  • DV2 Finetune: It pretrains a DreamerV2 agent [10] on distracting videos and then finetunes the trained model in the downstream tasks.
+Note that some tasks have different action spaces, which makes it difficult to finetune directly.
+Therefore, the action space of two tasks is set as the maximum action space of both environments.
+  • TED [7]: It adopts a classification task to learn temporally disentangled representations in visual RL.
+  • CURL [18]: A model-free RL method that employs contrastive learning to improve its sample efficiency.
+Appendix B Behavior Learning
+For the behavior learning of DisWM, we adopt the actor-critic method following DreamerV2 [10].
+Concretely, the actor and critic are both implemented as MLPs with ELU activations [5].
+Formally, the actor and critic are defined as below:
+Actor: $\hat{a}_{t}\sim\pi_{\psi}(\hat{a}_{t}|\hat{z}_{t})$ (6)
+Critic: $v_{\xi}(\hat{z}_{t})\approx\mathbb{E}_{p_{\phi},p_{\psi}}\Big{[}\textstyle\sum_{\tau\geq t}\hat{\gamma}_{\tau-t}\hat{r}_{\tau}\Big{]}.$
+The actor $\pi_{\psi}$ is optimized by maximizing
+$\mathcal{L}(\psi)=\ \mathbb{E}_{p_{\phi},p_{\psi}}\Big{[}\sum_{t=1}^{H-1}(\underbrace{\beta\mathrm{H}\left[a_{t}\mid\hat{z}_{t}\right]}_{\text{entropy regularization}}+\underbrace{\rho V_{t}}_{\text{dynamics backprop}}$ (7)
+$+\underbrace{(1-\rho)\ln\pi_{\psi}(\hat{a}_{t}\mid\hat{z}_{t})\texttt{sg}(V_{t}-v_{\xi}(\hat{z}_{t}))}_{\text{REINFORCE}}\Big{]}.$
+We train the critic $v_{\xi}$ by minimizing
+$\mathcal{L}(\xi)=\mathbb{E}_{p_{\phi},p_{\psi}}\Big{[}\sum_{t=1}^{H-1}\frac{1}{2}\left(v_{\xi}\left(\hat{z}_{t}\right)-\texttt{sg}\left(V_{t}\right)\right)^{2}\Big{]}.$ (8)
+where sg is a stop gradient operator.
+The $\lambda$-target $V_{t}$ that involves a weighted average of reward information used in Eq. (7) and Eq. (8) is defined as:
+$V_{t}\doteq\hat{r}_{t}+\hat{\gamma}_{t}\begin{cases}(1-\lambda)v_{\xi}\left(\hat{z}_{t+1}\right)+\lambda V_{t+1}&\text{ if }t<H\\ v_{\xi}\left(\hat{z}_{H}\right)&\text{ if }t=H\end{cases}.$ (9)
+where $H$ is the imagination horizon. Notably, the disentangled world model is not optimized during behavior learning.
+Appendix C Additional Results
+C.1 Results on DMC
+We compare the performance of DreamerV3 [12], TD-MPC2 [13], ContextWM [39], and our approach on DMC. As shown Table A, DisWM outperforms other strong baselines in terms of episode return.
+Model Reacher Easy $\rightarrow$ Cheetah Run Walker Walk $\rightarrow$ Humanoid Walk
+DreamerV3 662 $\pm$ 9 12 $\pm$ 17
+TD-MPC2 510 $\pm$ 15 1 $\pm$ 0
+ContextWM 661 $\pm$ 49 1 $\pm$ 0
+DisWM 817 $\pm$ 59 147 $\pm$ 85
+Table A: Comparison with strong baselines on DMC.
+C.2 Results on DrawerWorld
+We present results on DrawerWorld [37] in Table B. As reported in Table B, DisWM (source: Finger Spin) outperforms other baselines in terms of success rate (%) on all tasks.
+Model DrawerClose DrawerOpen
+TDMPC2 3 $\pm$ 6 43 $\pm$ 25
+ContextWM 37 $\pm$ 12 23 $\pm$ 25
+DisWM 77 $\pm$ 6 70 $\pm$ 10
+Table B: Performance on DrawerWorld with texture variations.
+C.3 Sensitivity of the Latent Space Dimension
+We visualize sensitivity analyses on the latent space dimension in Figure I.
+We observe that when $\mathbf{z}_{\text{dim}}$ for the $\beta$-VAE is too small, it impedes the learning of disentangled representations, leading to a decline in performance.
+Figure I: Sensitivity analyses on Cheetah Run $\rightarrow$ Walker Walk
+C.4 Runtime Comparisons
+We provide the detailed runtime and parameter comparisons with baselines in Table C. Note that the inference time is computed for one episode.
+Table C: Runtime and model size comparisons evaluated on DMC (Finger Spin $\rightarrow$ Reacher Easy). DV2 FT is short for DreamerV2 finetune.
+Model Training Steps Training time Inference time Params (M)
+CURL 100k 303 min 4.97 sec 10.7
+DV2 FT 200k 1522 min 9.88 sec 12.1
+APV 200k 1722 min 10.15 sec 13
+TED 100k 1051 min 20.49 sec 11.5
+DV2 100k 901 min 9.59 sec 12.1
+DisWM 200k 1311 min 9.48 sec 5.8
+C.5 Sample Diversity Visualization
+The adaptation stage enriches the sample diversity, as shown in Figure J, for Cheetah Run $\rightarrow$ Walker Walk, we sample 200 video clips of length 50 and visualize the corresponding latent features using t-SNE [25].
+We find that the latent features of the online interactions are more diverse than those of the offline dataset.
+Figure J: Sample diversity enhanced by adaptation.
+Appendix D Hyperparameters
+The final hyperparameters of DisWM are reported in Table D.
+Table D: Hyperparameters of DisWM.
+Name Notation Value
+Video prediction model
+Image size $64\times 64$
+KL divergence scale $\beta_{1}$ 11
+Disentanglement scale $\beta_{2}$ $0.015$
+Latent dimension 2020
+Learning rate $3\cdot 10^{-4}$
+Disentangled World Model
+Latent distillation weight $\eta$ $0.1$
+Disentanglement scale $\beta$ $0.015$
+KL divergence scale $\alpha$ 11
+Latent dimension 2020
+Batch size $B$ $50$
+Batch length $L$ $50$
+Learning rate $3\cdot 10^{-4}$
+Behavior Learning
+Imagination horizon $H$ $15$
+Discount $\gamma$ $0.99$
+$\lambda$-target $\lambda$ $0.95$
+Actor learning rate $8\cdot 10^{-5}$
+Critic learning rate $8\cdot 10^{-5}$